# Super-extensions of tensor algebras and their applications

### v1

@@ 7.1. Euler operators
$[k]=\frac{q^{k}-q^{-k}}{q-q^{-1}}\quad\mbox{and}\quad[k]!=[1][2]\cdots[k]$
@@
$[K_{i}:a]=\frac{q_{i}^{a}K_{i}-q_{i}^{-a}K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
$[K_{i}]^{h}_{!}=[K_{i}:0][K_{i}:-1]\cdots[K_{i}:1-h],\quad\mathrm{for}\ h\in\mathbb{N_{+}},$
and $[K_{i}]^{0}_{!}=1$. Then by Lemma 6.2 we have
@@
Set =$\mathcal{I}=\{(i_{d},i_{d-1},...,i_{1})\in[m+n]^{d}~{}|~{}i_{d}\leq i_{d-1}\leq\cdots\leq i_{1}\}$. For any $I\in\mathcal{I}$, let $[\mathbf{m}(I)]!=[\mathbf{m}_{1}(I)]![\mathbf{m}_{2}(I)]!\cdots[\mathbf{m}_{m+n}(I)]!$ where $\mathbf{m}_{k}(I)$ is the multiplicity of $k$ in $I$. By (7.1) and Proposition 6.1 we have
@@
Note that, for any $I,J\in\mathcal{I}$, we have
$[K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=\begin{cases}[\mathbf{m}_{i}(I)]!e_{J},&\text{if }\mathbf{m}_{i}(I)=\mathbf{m}_{i}(J);\\[3.0pt] 0,&\text{if }\mathbf{m}_{i}(I)>\mathbf{m}_{i}(J).\end{cases}$
Then by (7.2), one has
@@ 7.2. Descriptions of commutants
For any $d\geq 1$ and $0\leq k\leq d$, we denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
@@
For any $T\in\mathscr{H}_{d+k}$, we define $\pi_{d+k}(T)\in\mathrm{End}_{\mathbb{C}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ by $\pi_{d+k}(T)(\varphi)=\varphi\cdot T$ where $\varphi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$. Denote $\mathcal{B}_{d+k}=\pi_{d+k}(\mathscr{H}_{d+k})$.
@@
Asumme that $[d]!\neq 0$. We have
$\mathrm{Comm}(\mathcal{A}_{d}^{k})=\mathcal{B}_{d+k},$
@@
$\mathrm{Comm}(\mathcal{B}_{d+k})=\mathcal{A}_{d}^{k}.$
@@
It is easy to see that $\mathcal{A}_{d}^{k}\subset\mathrm{Comm}(\mathcal{B}_{d+k})$. Since $\mathcal{H}_{d}(q)$ is semisimple, by the dobule commutant theorem, it suffices to show $\mathrm{Comm}(\mathcal{B}_{d+k})\subset\mathcal{A}_{d}^{k}$.
For $f\in\mathrm{Comm}(\mathcal{B}_{d+k}),\phi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have
@@
where $\tau_{I}=L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\in\mathscr{H}_{d+k}$. So, $f=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})$. Since $L(f(e_{i_{d}}\cdots e_{i_{1}}))\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have $f\in\mathcal{A}_{d}^{k}$. ∎
Theorem 7.3 (Mitsuhashi).
Asumme that $[d]!\neq 0$. For each $d\geq 1$, we have
$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
@@
$\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))=\rho_{d}(\mathbf{U}_{m|n}),$
where $\rho_{d}$ denotes the representation of $\mathbf{U}_{m|n}$ on $\mathbb{V}^{\otimes d}\cong\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d}$.
@@
We just need to prove that $\rho_{d}(\mathbf{U}_{m|n})=\mathcal{A}_{d}^{0}$.
Proposition 6.8 shows that $\mathcal{A}_{d}^{0}\subset\rho_{d}(\mathbf{U}_{m|n})$. The other side is proved as follows. We first notice that $K_{i}\in\mathrm{Comm}(\pi_{d}(\mathscr{H}_{d}(q)))$. In the proof of Theorem 7.2, we have seent that $K_{i}\in\mathcal{A}_{d}^{0}$. Actually, we have
@@
Since Theorem 6.6 shows that $E_{ij}=L(e_{i})L(e_{j}^{*})K_{i}$ and $E_{ij}=K_{j}^{-1}L(e_{j})L(e_{i}^{*})$ for $i<j$ on ${\mathbb{V}^{\otimes d}}$, these elements commute with $\mathcal{B}_{d}$, then $\rho_{d}(\mathbf{U}_{m|n})\subset\mathcal{A}_{d}^{0}$. ∎

### v2

@@ 7.1. Euler operators
$k]=\frac{q^{k}-q^{-k}}{q-q^{-1}}\quad\mbox{and}\quad[k]!=[1][2]\cdots[k]$
@@
$K_{i}:a]=\frac{q_{i}^{a}K_{i}-q_{i}^{-a}K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
$K_{i}]^{h}_{!}=[K_{i}:0][K_{i}:-1]\cdots[K_{i}:1-h],\quad\mathrm{for}\ h\in\mathbb{N_{+}},$
and $K_{i}]^{0}_{!}=1$. Then by Lemma 6.2 we have
@@
Set $\mathcal{I}=\{(i_{d},i_{d-1},...,i_{1})\in[m+n]^{d}~|~i_{d}\leq i_{d-1}\leq\cdots\leq i_{1}\}$. For any $I\in\mathcal{I}$, let $\mathbf{m}(I)]!=[\mathbf{m}_{1}(I)]![\mathbf{m}_{2}(I)]!\cdots[\mathbf{m}_{m+n}(I)]!$ where $\mathbf{m}_{k}(I)$ is the multiplicity of $k$ in $I$. By (7.1) and Proposition 6.1 we have
@@
For any $I,J\in\mathcal{I}$, if $\mathbf{m}_{i}(I)=\mathbf{m}_{i}(J)$, we have
$K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=[\mathbf{m}_{i}(I)]!e_{J},$
and if $\mathbf{m}_{i}(I)>\mathbf{m}_{i}(J)$, we have
$K_{i}]^{\mathbf{m}_{i}(I)}_{!}.e_{J}=0.$
In the case where $\mathbf{m}_{i}(I)<\mathbf{m}_{i}(J)$ for some $i$, there must exist some index $j$ such that $\mathbf{m}_{j}(I)>\mathbf{m}_{j}(J)$. Thus by (7.2), one has
@@ 7.2. Descriptions of commutants
For any $d\geq 1$ and $0\leq k\leq d$, we denote by $\mathcal{A}_{d}^{k}$ the subspace of $\mathrm{End}_{\mathbb{C}(q)}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$ spanned by operators of the form
@@
The representation $\pi_{d+k}$ of $\mathscr{H}_{d+k}$ on $\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$ is given by right multiplication and we identify $\pi_{d+k}(\mathscr{H}_{d+k})$ with $\mathscr{H}_{d+k}$.
@@
Asumme that $d]!\neq 0$. We have
$\mathcal{A}_{d}^{k}=\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}),$
@@
$\mathrm{End}_{\mathcal{A}_{d}^{k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})=\mathscr{H}_{d+k}.$
@@
It is easy to see that $\mathcal{A}_{d}^{k}\subset\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})$. Since $\mathcal{H}_{d}(q)$ is semisimple, by the dobule commutant theorem, it suffices to show $\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k})\subset\mathcal{A}_{d}^{k}$.
For $f\in\mathrm{End}_{\mathscr{H}_{d+k}}(\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}),\phi\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have
@@
where $\tau_{I}=L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*})\phi\in\mathscr{H}_{d+k}$. So
$f=\sum_{I\in\mathcal{I}}\frac{1}{[\mathbf{m}(I)]!}L(f(e_{i_{d}}\cdots e_{i_{1}}))L(e_{i_{1}}^{*})\cdots L(e_{i_{d}}^{*}).$
Since $L(f(e_{i_{d}}\cdots e_{i_{1}}))\in\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d+k}$, we have $f\in\mathcal{A}_{d}^{k}$. ∎
Corollary 7.3 (Mitsuhashi).
Asumme that $d]!\neq 0$. For each $d\geq 1$, we have
$\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))=\mathrm{End}_{\mathscr{H}_{d}}(\mathbb{V}^{\otimes d}),$
@@
$\mathrm{End}_{\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))}(\mathbb{V}^{\otimes d})=\mathscr{H}_{d},$
where $\rho_{d}$ denotes the representation of $U_{q}(\mathfrak{gl}(m|n))$ on $\mathbb{V}^{\otimes d}\cong\mathbb{V}^{\otimes d}\otimes_{\mathscr{H}_{d}}\mathscr{H}_{d}$.
@@
We just need to prove that $\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))=\mathcal{A}_{d}^{0}$.
Proposition 6.8 shows that $\mathcal{A}_{d}^{0}\subset\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))$. The other side is proved as follows. We first notice that $K_{i}\in\mathrm{End}_{\mathscr{H}_{d}}(\mathbb{V}^{\otimes d})$. In the proof of Theorem 7.2, we have seen that $K_{i}\in\mathcal{A}_{d}^{0}$. Actually, we have
@@
It follows from Theorem 6.6 that $E_{ij}=L(e_{i})L(e_{j}^{*})K_{i}$ and $E_{ij}=K_{j}^{-1}L(e_{j})L(e_{i}^{*})$ for $i<j$ on $\mathbb{V}^{\otimes d}$}. So these elements commute with $\mathcal{B}_{d}$, and hence $\rho_{d}(U_{q}(\mathfrak{gl}(m|n)))\subset\mathcal{A}_{d}^{0}$. ∎